# Distributionally Robust Active Learning for Gaussian Process Regression

### v1

@@ 1 Introduction
Active learning (AL) (Settles, 2009) is a framework for achieving high prediction performance with fewer data when labeling new data is expensive. For this purpose, AL algorithms actively acquire the label of data that improves the prediction performance of some statistical model based on acquisition functions (AFs). Many types of AFs have been proposed, such as uncertainty sampling (US), random sampling (RS), variance reduction, and information gain, as summarized in Settles (2009).
Gaussian process regression (GPR) model (Rasmussen and Williams, 2005) is often used as a base statistical model for AL algorithms due to its flexible prediction capability (Seo et al., 2000; Yu et al., 2006; Guestrin et al., 2005; Krause et al., 2008b; Hoang et al., 2014). Standard AL methods for the GPR are based on information gaion (Guestrin et al., 2005; Krause et al., 2008b; Kirsch et al., 2021; Kirsch and Gal, 2022; Bickford Smith et al., 2023). Most information gain-based approaches are heuristics without theoretical guarantees. A notable exception is the work by Guestrin et al. (2005); Krause et al. (2008b), which shows the US for the GPR model is optimal to maximize the information gain from the obtained data labels regarding GP prior. Furthermore, from the analysis of kernelized bandits (e.g., Srinivas et al., 2010; Salgia et al., 2024), we can see that the US and RS guarantee the convergence of the maximum of posterior variance (See Proposition 2.3 for details). Another commonly used AF is variance reduction (Seo et al., 2000; Yu et al., 2006; Shoham and Avron, 2023), which can be computed efficiently in the GPR. However, these AFs do not incorporate the importance of the unlabelled dataset, that is, the prior information regarding the target distribution. In addition, to our knowledge, except for the worst-case analysis in Proposition 2.3, there are no theoretical guarantees for the target prediction error.
@@
This paper develops a DRAL framework for the GPR model. We aim to minimize the worst-case expected error, where the worst-case scenario and the expectation are taken regarding the target distribution candidates and chosen target distributions, respectively. Note that our formulation is a generalization of target distribution-aware AL since it includes the case in which the unique target distribution can be specified. We perform the theoretical analysis under two conditions called Bayesian and frequentist assumptions (Srinivas et al., 2010), in which we leverage several useful lemmas in kernelized bandit literature (Srinivas et al., 2010; Vakili et al., 2021a, b; Kusakawa et al., 2022).
@@ 2.1 GPR model
We will use the following useful result from Theorem E.4 in Kusakawa et al. (2022):
@@
2.2 Uncerrainty Sampling and Random Sampling
@@
For the US, see, e.g., Eq. (16) in Vakili et al. (2021a) and Lemma 5.4 in Srinivas et al. (2010). For the RS, see Theorem 3.1 in Salgia et al. (2024). ∎
@@ 2.3.1 Bayesian Assumption
As with Lemma 5.1 in Srinivas et al. (2010), the credible interval can be obtained as follows:
@@ 2.3.2 Frequentist Assumption
We assume that $f$ is an element of the reproducing kernel Hilbert space (RKHS) specified by the kernel $k$ as with Srinivas et al. (2010); Chowdhury and Gopalan (2017); Vakili et al. (2021a, 2022); Li and Scarlett (2022):
@@
Furthermore, for continuous ${\cal X}$, we assume the following smoothness condition as with Chowdhury and Gopalan (2017); Vakili et al. (2021a, 2022):
@@
Lemma 2.9 (Lemma 5.1 in De Freitas et al. (2012)).
@@
We rely on the confidence bounds for non-adaptive sampling methods, which is a direct consequence of Theorem 1 in Vakili et al. (2021a) and the union bound:
@@
3 Problem Statements and Its Property
This section provides our problem setup and its properties.
@@ 3.2 High Probability Bound of Error
Furthermore, we need the Lipschitz constant of $\mu_{T}$. In the frequentist setting, the Lipschitz constant for $\mu_{T}$ can be derived as ${\cal O}(L_{k}\sqrt{t\log t})$ by Lemma 4 in Vakili et al. (2021a) and Lemma 2.9. To obtain a slightly tighter upper bound, we show the following lemma:
@@ 4 Proposed Methods and Analysis
We aim to design algorithms that enjoy both a similar convergence guarantee as the US and RS and practical effectiveness incorporating the information of ${\cal P}$. In particular, we consider two algorithms inspired by the greedy algorithm and the RS and show theoretical guarantees. Algorithm 1 shows the pseudo-code of proposed algorithms.
@@ 4.1 Algorithms
Second, we consider the greedy algorithm since its practical efficiency has often been reported (e.g., Bian et al., 2017). However, in our setup, the greedy algorithm should be
@@ 4.2 Analysis
See Appendix B for the proof, in which Lemma 3 in Kirschner and Krause (2018) is used to show Theorem 4.1.
@@ 5 Related Work
In OED or AL, subset selection algorithms (Das and Kempe, 2008) are often leveraged. The subset selection is a general problem whose goal is finding the subset that maximizes some set function. Therefore, the AL can be seen as the subset selection of $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{t}\in{\cal X}$. In this literature, the submodular property of the set function, for which the greedy algorithm can be optimal, is commonly investigated (Das and Kempe, 2008; Krause et al., 2008b; Guestrin et al., 2005; Bian et al., 2017). The criteria for the AL, such as the D-optimality of the GPR (Krause et al., 2008b; Guestrin et al., 2005), sometimes satisfy the submodular property. Furthermore, Das and Kempe (2008) have shown sufficient conditions for that the greedy algorithms are optimal in Theorem 3.4 (an assumption can be rephrased as $k(\boldsymbol{x},\boldsymbol{x}^{\prime})\leq\frac{1}{4T}$ in our problem) and Section 8. However, even if we consider minimizing $\mathbb{E}_{p(\boldsymbol{x}^{*})}\left[\sigma_{T}^{2}(\boldsymbol{x}^{*})\right]$ with $|{\cal P}|=1$, these conditions and the submodularity do not hold in general. Therefore, the DR maximization of submodular function (e.g., Krause et al., 2008a; Staib et al., 2019) also cannot be applied directly.
Several studies have discussed the target distribution-aware AL. At least from Sugiyama (2005), the effectiveness of AL incorporating the information of target distribution for misspecified models has been discussed. Transductive AL (Seo et al., 2000; Yu et al., 2006; Shoham and Avron, 2023) can be interpreted as the expected error minimization when the uniformly random target distribution $p(\boldsymbol{x})=1/|{\cal X}|$ is specified. Kirsch et al. (2021); Kirsch and Gal (2022); Bickford Smith et al. (2023) extended this setting so that an arbitrary target distribution can be considered. For the non-Bayesian model, Frogner et al. (2021) further extended to DRAL using the AF called expected model change (Settles, 2009). However, most existing methods are heuristic greedy algorithms and are not theoretically guaranteed for the prediction error.
The DRAL is inspired by the DR learning (DRL) (Chen and Paschalidis, 2018; Chen et al., 2020). DRL considers learning the robust statistical model by optimizing the model parameter so that the worst-case expected loss is minimized, where the worst-case is taken regarding the target distribution candidates called an ambiguity set. Therefore, DRAL is an intuitive extension of DRL to AL.
@@
In addition, level set estimation (LSE) (Gotovos et al., 2013; Bogunovic et al., 2016; Inatsu et al., 2024) is an AL framework using the GPR model, which aims to classify the test input set by whether or not a black-box function value proceeds a given threshold. In particular, Inatsu et al. (2021) considers the variant of LSE, which aims to classify by whether or not the DR measure defined by the black-box function proceeds a given threshold. Thus, our problem setup differs from the problem of Inatsu et al. (2021).
Figure 2: Result of the error $E_{T}$ in the real-world data experiments with $\eta=0,0.001,0.01,0.1$. The horizontal and vertical axes show the number of iterations and $E_{T}$, respectively. The error bar shows mean and standard errors for 20 random trials regarding the random initial point (and the algorithm’s randomness). The top, middle, and bottom columns represent the results of the King County house sales dataset, red wine quality dataset, and auto MPG dataset, respectively.
@@ 6 Experiments
In this section, we demonstrate the effectiveness of the proposed methods via synthetic and real-world datasets. We employ RS, US, variance reduction (Yu et al., 2006), and expected predictive information gain (EPIG) (Bickford Smith et al., 2023) as the baseline. We show the implementation details of EPIG in Appendix C.2. Furthermore, as the ablation study, we performed the method, referred to as DR variance reduction, that greedily minimizes $E_{p_{t}(\boldsymbol{x}^{*})}\left[\sigma^{2}_{t}(\boldsymbol{x}^{*}\mid\boldsymbol{x})\right]$, that is, the unconstrained version of Eq. (4). We referred to the proposed methods as DR random and constrained DR variance reduction (CDR variance reduction). We evaluate the performance by the error $E_{T}$ defined in Eq. (2). Furthermore, for the synthetic dataset, we show the result of $\max_{p\in{\cal P}}\mathbb{E}_{p(\boldsymbol{x})}\left[\sigma_{t}^{2}(\boldsymbol{x})\mid{\cal D}_{t}\right]$ in Appendix C.1.
@@ 6.1 Synthetic Data Experiemnts
We set ${\cal X}=\{-1,-0.8,\ldots,1\}^{3}$, where $|{\cal X}|=11^{3}=1331$. The target function $f$ is the sample path from GPs, where we use SE and Matérn-$\nu$ kernels with $\nu=5/2$. We use the fixed hyperparameters of the kernel function in the GPR model, which is used to generate $f$, and fix $\sigma^{2}=10^{-4}$. The first input $\boldsymbol{x}_{1}$ is selected by uniformly random, and $T$ is set as $400$. Furthermore, we set $p_{\rm ref}={\cal N}(\boldsymbol{0},0.2\boldsymbol{I}_{3})$.
Figures 1 shows the result. We can see that DR and CDR variance reductions show superior performance consistently for all the kernel functions and $\eta$, although the DR random is often inferior to those due to the randomness. This result suggests that the DR and CDR variance reductions effectively incorporate the information of ${\cal P}$. Furthermore, although the constraint by ${\cal X}_{t}$ is required for the theoretical analysis in CDR variance reduction, we can confirm that it does not sacrifice the practical effectiveness. On the other hand, the usual AL methods, such as US and variance reduction, deteriorate when $\eta$ is small since they do not incorporate the information of ${\cal P}$. When $\eta$ is large since our problem approaches the worst-case error minimization, the US and variance reduction result in relatively good results. On the other hand, the EPIG designed for the case $\eta=0$ is inferior for all $\eta$ since the EPIG is based on the entropy ${\cal O}\left(\log(\sigma_{t}(\boldsymbol{x}))\right)$, not the squared error. That is, since $\frac{\partial\log a}{\partial a}=\frac{1}{a}$ is large when $a\rightarrow 0$, the EPIG tends to make a small $\sigma_{t}(\boldsymbol{x})$ more small than making a large $\sigma_{t}(\boldsymbol{x})$ small.

### v2

@@ 1 Introduction
Active learning (AL) (Settles, 2009) is a framework for achieving high prediction performance with fewer data when labeling new data is expensive. For this purpose, AL algorithms actively acquire the label of data that improves the prediction performance of some statistical model based on acquisition functions (AFs). Many types of AFs have been proposed, such as uncertainty sampling (US), random sampling (RS), variance reduction, and information gain, as summarized in (Settles, 2009).
Gaussian process regression (GPR) model (Rasmussen and Williams, 2005) is often used as a base statistical model for AL algorithms due to its flexible prediction capability (Seo et al., 2000; Yu et al., 2006; Guestrin et al., 2005; Krause et al., 2008b; Hoang et al., 2014; Hübotter et al., 2024). Standard AL methods for the GPR are based on information gain (Guestrin et al., 2005; Krause et al., 2008b; Kirsch et al., 2021; Kirsch and Gal, 2022; Bickford Smith et al., 2023; Hübotter et al., 2024). Most information gain-based approaches are heuristics without theoretical guarantees. A notable exception is the work by (Guestrin et al., 2005; Krause et al., 2008b), which shows that the US for the GPR model is optimal to maximize the information gain from the obtained data labels regarding the GP prior. Furthermore, from the analysis of kernelized bandits (e.g., Srinivas et al., 2010; Salgia et al., 2024), we can see that the US and RS guarantee the convergence of the maximum of posterior variance (See Proposition 2.3 for details). Another commonly used AF is variance reduction (Seo et al., 2000; Yu et al., 2006; Shoham and Avron, 2023; Hübotter et al., 2024), which can be computed efficiently in the GPR. However, these AFs do not incorporate the importance of the unlabelled dataset, that is, the prior information regarding the target distribution. In addition, to our knowledge, except for the worst-case analysis in Proposition 2.3, there are no theoretical guarantees for the target prediction error.
@@
This paper develops a DRAL framework for the GPR model. We aim to minimize the worst-case expected error, where the worst-case scenario and the expectation are taken regarding the target distribution candidates and chosen target distributions, respectively. Note that our formulation generalizes target distribution-aware AL since it includes the case in which the unique target distribution can be specified. We perform the theoretical analysis under two conditions called Bayesian and frequentist assumptions (Srinivas et al., 2010), in which we leverage several useful lemmas in kernelized bandit literature (Srinivas et al., 2010; Vakili et al., 2021a, b; Kusakawa et al., 2022).
@@ 2.1 GPR model
We will use the following useful result from Theorem E.4 in (Kusakawa et al., 2022):
@@
2.2 Uncertainty Sampling and Random Sampling
@@
For the US, see, e.g., Eq. (16) in (Vakili et al., 2021a) and Lemma 5.4 in (Srinivas et al., 2010). For the RS, see Theorem 3.1 in (Salgia et al., 2024). ∎
@@ 2.3.1 Bayesian Assumption
As with Lemma 5.1 in (Srinivas et al., 2010), the credible interval can be obtained as follows:
@@ 2.3.2 Frequentist Assumption
We assume that $f$ is an element of the reproducing kernel Hilbert space (RKHS) specified by the kernel $k$ as with (Srinivas et al., 2010; Chowdhury and Gopalan, 2017; Vakili et al., 2021a, 2022; Li and Scarlett, 2022):
@@
Furthermore, for continuous ${\cal X}$, we assume the following smoothness condition as with (Chowdhury and Gopalan, 2017; Vakili et al., 2021a, 2022):
@@
Lemma 2.9 (Lemma 5.1 in (De Freitas et al., 2012)).
@@
We rely on the confidence bounds for non-adaptive sampling methods, which is a direct consequence of Theorem 1 in (Vakili et al., 2021a) and the union bound:
@@
3 Problem Statement and Its Property
This section provides details on our problem setup and its properties.
@@ 3.2 High Probability Bound of Error
Furthermore, we need the Lipschitz constant of $\mu_{T}$. In the frequentist setting, the Lipschitz constant for $\mu_{T}$ can be derived as ${\cal O}(L_{k}\sqrt{t\log t})$ by Lemma 4 in (Vakili et al., 2021a) and Lemma 2.9. To obtain a slightly tighter upper bound, we show the following lemma:
@@ 4 Proposed Methods and Analysis
We aim to design algorithms that enjoy both a similar convergence guarantee as the US and RS and practical effectiveness, incorporating the information of ${\cal P}$. In particular, we consider two algorithms inspired by the greedy algorithm and the RS and show theoretical guarantees. Algorithm 1 shows the pseudo-code of the proposed algorithms.
@@ 4.1 Algorithms
Second, we consider the greedy algorithm since its practical efficiency has often been reported (e.g., Bian et al., 2017). However, in our setup, the algorithm that greedily decreases the expected posterior variance should be
@@
Necessity of Constraints:
We considered that the constraint regarding $\sigma^{2}_{t-1}(\boldsymbol{x})$ makes the analysis easy since the US that maximize $\sigma_{t-1}(\boldsymbol{x}_{t})$ achieves the error convergence, as shown in Proposition 2.3. Therefore, we employ the constraint on ${\cal X}_{t}$. We set the threshold of the constraint as $\mathbb{E}_{p_{t}(\boldsymbol{x}^{*})}[\sigma_{t-1}^{2}(\boldsymbol{x}^{*})]$ sake of the analysis. On the other hand, our experimental results suggest that the greedy (approximated) expected error reduction algorithm without the constraint shows superior performance. Therefore, removing or alleviating the constraint can be important future work.
Computational Complexity:
The computation of the GP has $O(T^{3})$ computational complexity, which can be alleviated by scalable GP learning approaches (Liu et al., 2020). However, more careful proofs incorporating an approximation error in GP learning are required to derive similar theoretical analyses as ours, as with (Vakili et al., 2022). On the other hand, the computational complexity of the maximization $\max_{p\in{\cal P}}\mathbb{E}_{p(\boldsymbol{x}^{*})}[\sigma^{2}_{t-1}(\boldsymbol{x}^{*})]$, the expectation $\mathbb{E}_{p_{t}(\boldsymbol{x}^{*})}[\sigma^{2}_{t-1}(\boldsymbol{x}^{*})]$ and the sampling from $p_{t}(\boldsymbol{x})$ depends on the ambiguity set ${\cal P}$ and may increase in proportion to $d$. Although our experiments focus on the set of discrete probability distributions with a moderate size of $|{\cal X}|$, the above computations may be complicated if ${\cal P}$ is the set of continuous probability distributions or $|{\cal X}|$ is huge. Extensions to such more computationally intractable ambiguity sets ${\cal P}$, e.g., the ball defined by Wasserstein distance and Kullback–Leibler divergence (Hu and Hong, 2013; Frogner et al., 2021), is crucial future work.
@@ 4.2 Analysis
See Appendix B for the proof, in which Lemma 3 in (Kirschner and Krause, 2018) is used to show Theorem 4.1.
@@ 5 Related Work
In OED or AL, subset selection algorithms (Das and Kempe, 2008) are often leveraged. The subset selection is a general problem whose goal is to find the subset that maximizes some set function. Therefore, the AL can be seen as the subset selection of $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{t}\in{\cal X}$. In this literature, the submodular property of the set function, for which the greedy algorithm can be optimal, is commonly investigated (Das and Kempe, 2008; Krause et al., 2008b; Guestrin et al., 2005; Bian et al., 2017). The criteria for the AL, such as the D-optimality of the GPR (Krause et al., 2008b; Guestrin et al., 2005), sometimes satisfy the submodular property. Furthermore, Das and Kempe (2008) have shown sufficient conditions for the greedy algorithms to be optimal in Theorem 3.4 (an assumption can be rephrased as $k(\boldsymbol{x},\boldsymbol{x}^{\prime})\leq\frac{1}{4T}$ in our problem) and Section 8. However, even if we consider minimizing $\mathbb{E}_{p(\boldsymbol{x}^{*})}\left[\sigma_{T}^{2}(\boldsymbol{x}^{*})\right]$ with $|{\cal P}|=1$, these conditions and the submodularity do not hold in general. Therefore, the DR maximization of submodular function (e.g., Krause et al., 2008a; Staib et al., 2019) also cannot be applied directly.
Several studies have discussed the target distribution-aware AL. At least from (Sugiyama, 2005), the effectiveness of AL incorporating the information of target distribution for misspecified models has been discussed. Transductive AL (Seo et al., 2000; Yu et al., 2006; Shoham and Avron, 2023) can be interpreted as the expected error minimization when the uniformly random target distribution $p(\boldsymbol{x})=1/|{\cal X}|$ is specified. Kirsch et al. (2021); Kirsch and Gal (2022); Bickford Smith et al. (2023) extended this setting so that an arbitrary target distribution can be considered. These existing methods are heuristic algorithms that do not guarantee the convergence of the error. On the other hand, Hübotter et al. (2024) show transductive AL methods with theoretical guarantees. However, for the VTR algorithm in (Hübotter et al., 2024), their analysis assumes an assumption of submodularity, which may not hold as discussed immediately after Assumption 3.2 of (Hübotter et al., 2024). Furthermore, their analysis for the VTR algorithm results in $\max_{\boldsymbol{x}\in{\cal X}}\sigma^{2}_{T}(\boldsymbol{x})=O(|{\cal X}|\gamma_{T}/T)$, which is vacuous in the usual regime that $T<|{\cal X}|$ (see Theorems 3.3 and C.13 and Section C.6.2 of Hübotter et al. (2024)).
Frogner et al. (2021) further extended to DRAL using the AF called expected model change (Settles, 2009) for the non-Bayesian model. In addition, Liu et al. (2015) considers DRAL for non-Bayesian classification models. However, these methods are heuristic greedy algorithms and are not theoretically guaranteed for the prediction error.
The DRAL is inspired by the DR learning (DRL) (Chen and Paschalidis, 2018; Chen et al., 2020). DRL considers learning a robust statistical model by optimizing the model parameter so that the worst-case expected loss is minimized, where the worst-case is taken regarding the target distribution candidates called an ambiguity set. Therefore, DRAL is an intuitive extension of DRL to AL.
@@
In addition, level set estimation (LSE) (Gotovos et al., 2013; Bogunovic et al., 2016; Inatsu et al., 2024) is an AL framework using the GPR model, which aims to classify the test input set by whether or not a black-box function value exceeds a given threshold. In particular, Inatsu et al. (2021) consider the variant of LSE, which aims to classify by whether or not the DR measure defined by the black-box function exceeds a given threshold. Thus, our problem setup differs from the problem of (Inatsu et al., 2021).
Figure 2: Result of the expected squared error $E_{T}$ in the real-world data experiments with $\eta=0,0.001,0.01,0.1$. The horizontal and vertical axes show the number of iterations and $E_{T}$, respectively. The error bar shows mean and standard errors for 20 random trials regarding the random initial point (and the algorithm’s randomness). The top, middle, and bottom rows represent the results of the King County house sales dataset, red wine quality dataset, and auto MPG dataset, respectively.
@@ 6 Experiments
In this section, we demonstrate the effectiveness of the proposed methods via synthetic and real-world datasets. We employ RS, US, variance reduction (Yu et al., 2006), and expected predictive information gain (EPIG) (Bickford Smith et al., 2023) as the baseline. Note that we do not employ the method of (Frogner et al., 2021) since adapting their method to GPR models is not apparent, and they focus on the ambiguity sets defined by Wasserstein distance over a continuous input domain. We show the implementation details of EPIG in Appendix C.2. Furthermore, as the ablation study, we performed the method, referred to as DR variance reduction, that greedily minimizes $E_{p_{t}(\boldsymbol{x}^{*})}\left[\sigma^{2}_{t}(\boldsymbol{x}^{*}\mid\boldsymbol{x})\right]$, that is, the unconstrained version of Eq. (4). We referred to the proposed methods as DR random and constrained DR variance reduction (CDR variance reduction). We evaluate the performance by the error $E_{T}$ defined in Eq. (2). Furthermore, for the synthetic dataset, we show the result of $\max_{p\in{\cal P}}\mathbb{E}_{p(\boldsymbol{x})}\left[\sigma_{t}^{2}(\boldsymbol{x})\mid{\cal D}_{t}\right]$ in Appendix C.1.
@@
The aim of our experiments is to show that the proposed algorithms are consistently well-performing for any size of ${\cal P}$, i.e., $\eta$, in contrast to the baselines, which can deteriorate for some $\eta$. The same parameter $\eta$ is used for the computation of the performance measure $E_{T}$ and the actual algorithm. Therefore, we expect that our proposed algorithms show good performance among all $\eta=0,0.001,0.01,0.1$. On the other hand, if $\eta$ is large, then ${\cal P}$ starts to include more distributions. Therefore, the proposed algorithms behave similarly to the US when $\eta$ is large. Hence, the proposed algorithms have comparable performance to the usual AL methods when $\eta$ is large, though the proposed algorithms are superior to those if $\eta$ is small. On the other hand, EPIG is designed for the case of $\eta=0$, that is, the test distribution is explicitly specified. However, EPIG aims to decrease the entropy $\mathbb{E}_{p(\boldsymbol{x})}[\log(\sigma_{t-1}(\boldsymbol{x}))]$, not $\mathbb{E}_{p(\boldsymbol{x})}[\sigma^{2}_{t-1}(\boldsymbol{x})]$. Therefore, EPIG is not suitable for decreasing $E_{T}$ and can be inefficient in our experiments even for the case of $\eta=0$.
@@ 6.1 Synthetic Data Experiemnts
We set ${\cal X}=\{-1,-0.8,\ldots,1\}^{3}$, where $|{\cal X}|=11^{3}=1331$. The target function $f$ is the sample path from GPs, where we use SE and Matérn-$\nu$ kernels with $\nu=5/2$. We use the fixed hyperparameters of the kernel function in the GPR model, which is used to generate $f$, and fix $\sigma^{2}=10^{-4}$. The first input $\boldsymbol{x}_{1}$ is selected uniformly at random, and $T$ is set to $400$. Furthermore, we set $p_{\rm ref}={\cal N}(\boldsymbol{0},0.2\boldsymbol{I}_{3})$.
Figure 1 shows the result. We can see that DR and CDR variance reductions show superior performance consistently for all the kernel functions and $\eta$, although the DR random is often inferior to those due to the randomness. This result suggests that the DR and CDR variance reductions effectively incorporate the information of ${\cal P}$. Furthermore, although the constraint by ${\cal X}_{t}$ is required for the theoretical analysis in CDR variance reduction, we can confirm that it does not sacrifice the practical effectiveness. On the other hand, the usual AL methods, such as US and variance reduction, deteriorate when $\eta$ is small since they do not incorporate the information of ${\cal P}$. When $\eta$ is large, since our problem approaches the worst-case error minimization, the US and variance reduction result in relatively good results. On the other hand, the EPIG designed for the case $\eta=0$ is inferior for all $\eta$ since the EPIG is based on the entropy ${\cal O}\left(\log(\sigma_{t}(\boldsymbol{x}))\right)$, not the squared error.